Without her contact lenses, Marissa cannot clearly see an object more than 40.0 cm away.

(a) Calculate the focal length of the contact lenses.

(b) What power and type (diverging or converging) of lens should her contact lenses have to give her normal vision?

(c) Suppose Marissa's near point (without her contact lenses) is 15.0 cm. What is the closest object she can see clearly with her contact lenses on?


1
Expert's answer
2021-07-08T13:28:03-0400

Gives

u=-40cm

v=infinite

"\frac{1}{f}=\frac{1}{u}+\frac{1}{v}"

"\frac{1}{f}=\frac{1}{infinite}-\frac{1}{-40}"

"\frac{1}{f}=0+\frac{1}{40}"

"\frac{1}{f}=\frac{1}{40}"

"f=40cm"

f=0.4m

Power

"P=\frac{1}{f}=\frac{1}{0.4}=2.5D"


f=15cm

"P=\frac{1}{f}=\frac{1}{0.15}=6.67D"

closest object she can see clearly with her contact lenses P=6.67D